OpenStudy (mathshelp0201):

In a sale, the price of all jackets is decreased by 20% of the normal price. John buys a leather jacket with a normal price of £108. a) Work out the sale price of John's jacket. Ben buys a tweed jacket at a sale price of £96. b) Work out the normal price of the tweed jacket.

sam (.sam.):

So.. what is 20% of 108?

OpenStudy (mathshelp0201):

21.6

sam (.sam.):

Right, now 108 minus 21.6 will be the price of John's jacket when its on sale

OpenStudy (mathshelp0201):

so 86.4? I thought you had to do 108/1.2

sam (.sam.):

Why 1.2? You don't have to, you've worked out the discount price already. 21.6 isn't it

sam (.sam.):

In short, the discount price of 20% will be \[\frac{20}{100} \times 108 =£ 21.6\] Then the 'discounted' price from the original price will be \[100-21.6=£78.4\]

sam (.sam.):

'108'-21.6=£86.4 Sorry